# Open Quantum System Approaches to Superconducting Qubits

### v1

@@ Abstract
Superconducting qubits are currently a leading platform for developing a scalable quantum computer. However, random and uncontrollable noises from the environment during the design and measurement of qubits lead to limitations in qubit coherence time and gate fidelity, which is a major challenge in current state-of-art for superconducting quantum computing. To advance superconducting qubit technologies, it is essential to understand and mitigate environmentally-induced errors. This requires modeling superconducting qubits as open quantum systems coupled to their surroundings. The present study aims to provide useful open quantum system approaches to analyze and quantify the interaction between superconducting qubits and their environments. We provide an accessible introduction to open quantum system formalism for newcomers to the field. For experts, we discuss recently developed methods for analyzing superconducting qubit dynamics under realistic noise. We outline how these techniques provide quantitative insights into decoherence mechanisms and how they can guide design improvements to enhance qubit coherence time. This self-contained review of open quantum system approaches can be used to model, understand, and improve superconducting qubit performance in the presence of unavoidable environmental noises.
Keywords decoherence, superconducting qubit, noise, circuit QED
@@ III.1 Basic Definitions
[…]
@@
in which the ground states of the qubit are eigenstates of charge. It means that they are only dependent on the number of excess Cooper pairs in a disconnected superconducting island. Single CPB and transmon are well-known examples.
@@
consists of a superconducting loop interrupted by one or more Josephson junctions. Their states match the direction in which the superconducting current flows in its loop. Unlike charge qubits, flux qubits have high coherence and a large anharmonicity. C-shunted flux qubits and fluxonium qubits are examples.
@@
which have bigger Josephson junctions compared to charge qubits. Their states relate to two levels of current energy in a Josephson junction which two levels are defined by quantum oscillations of the phase difference between the two sides of the junction. In phase qubits, the different states are usually determined by the direction of a flowing current in the Josephson loop.
@@ III.2 Hamiltonian
[…]
@@ III.3.1 Formalism of decoherence
The formalism of decoherence in superconducting qubits is similar to other quantum systems and includes two main steps: relaxation and dephasing. These mechanisms lead to the exponential vanishing of off-diagonal elements of the density matrix $\rho$ of the qubit which shows the overlap between two states of a qubit.
@@ III.3.2 Source of decoherence
is a common noise in solid-state devices. This noise originates fluctuating charges in the structural disorder or charge traps in the dielectrics and tunnel barrier of the junction. It can be modeled as an ensemble of fluctuating two-level systems or as bulk dielectric loss Wang ; Dial . Charge noise is mainly responsible for the relaxation ($T_{1}$) of a transmon qubit Krantz . To mitigate the effects of charge noise, we can use the etching and passivation to remove amorphous layers, crystalline dielectrics and ordered Josephson junction barriers Siddiqi .
@@ IV CQED
CQED is a new research field that was developed about twenty years ago Blais2 ; Wallraff ; Girvin . CQED uses theoretical and experimental methods to manipulate quantum states for technological applications. Currently, it plays an essential role in quantum information processing based on superconducting qubits Blais . Moreover, CQED is now a leading strategy in single photon detectors Besse ; Kono ; Lescanne , development of quantum-limited amplifiers Clerk1 ; Roy , hybrid quantum systems Clerk2 and semiconducting quantum dots Burkard .
@@
$H=\hbar\omega_{\rm c}(a^{\dagger}a+\frac{1}{2})+\frac{\hbar\omega_{\rm q}}{2}\sigma_{z}+\hbar g(a^{\dagger}\sigma^{-}+\sigma^{+}a)+H_{\rm c\varepsilon}+H_{\rm q\varepsilon},$ (6)
The first term represents cavity harmonic oscillator energy that $\omega_{\rm c}$ is the angular oscillation frequency of the relevant cavity mode. The second term shows the qubit energy in which the qubit frequency $\omega_{\rm q}$ is the energy difference between the ground state $|g\rangle$ and the excited state $|e\rangle$ of the qubit. The third term denotes the coupling between qubit and cavity with the strength of interaction $g$ in which the rotating wave approximation (RWA) is applied Knight ; Walls . This term shows the coherent exchange of excitation between the qubit and cavity and describes a dipole interaction where an atom can absorb ($\sigma^{+}a$) and emit ($a^{\dagger}\sigma^{-}$) a photon from/to the field at rate $g$. The two last terms show the coupling of cavity and qubit with the environment, respectively, that causes relaxation and dephasing of the qubit. Moreover, $a$ and $a^{\dagger}$ are annihilation and creation operators of the resonator photons, respectively and $\sigma_{z}=|e\rangle\langle g|-|g\rangle\langle e|$ is the atomic inversion operator and the operators $\sigma^{-}=|g\rangle\langle e|$ and $\sigma^{+}=|e\rangle\langle g|$ are the raising and lowering operators of the qubit. The detuning of a qubit from the cavity can be defined as $\omega_{\rm cq}=\omega_{\rm q}-\omega_{\rm c}$.
@@ V.1.1 Non-unitary evolution
To describe the state of an open quantum system as a mixed state, we use the density matrix formalism. A density matrix $\rho$ is a Hermitian operator describes a probability distribution of quantum states $\left|\psi_{n}\right>$ in a matrix representation that defined as $\rho=\sum_{n}p_{n}\left|\psi_{n}\right>\left<\psi_{n}\right|$, where $p_{n}$ is the probability that the system is in $\left|\psi_{n}\right>$. The normalization condition of the density matrix is defined as $\operatorname{Tr}\rho=1$ which ”$\operatorname{Tr}$” denotes Trace operation. Moreover, for a pure quantum state, we have $\operatorname{Tr}\rho^{2}=1$ and for a mixed state $\operatorname{Tr}\rho^{2}<1$. In general, the time evolution of a density matrix $\rho$ is the basis of the master equations in the theory of open quantum systems Schlosshauer1 ; Breuer .
@@ V.1.2 Derivation of the Lindblad master equation
$\frac{\operatorname{d}\rho_{\rm q}(t)}{\operatorname{d}t}=-i[H_{\rm q}(t),\rho_{\rm q}(t)]+\Gamma_{1}\Big{(}\sigma_{-}\rho_{\rm q}\sigma_{+}-\frac{1}{2}\{\sigma_{-}\sigma_{+},\rho_{\rm q}\}\Big{)}+\Gamma_{\phi}\Big{(}\sigma_{z}\rho_{\rm q}\sigma_{z}-\rho_{\rm q}\Big{)},$ (15)
@@ V.2.1 Framework
The evolution of a system and its environment is taken from the total Hamiltonian and traces over the degrees of freedom of the environment. Although various methods have been introduced in the approach of perturbative analysis of the interaction between the system and the environment such as Bloch-Redfield Bloch ; Redfield , Nakajima-Zwanzig Nakajima ; Zwanzig , Schoeller-Schon Schoeller , but finally all these methods reduce to Bloch-Redfield method.
@@
Recently, the Bloch-Redfield master equation has been used extensively to study the dissipative behavior of the qubits interacting with the environment Whitney . In the following, we analyze the formalism to derive the Bloch-Redfield master equation for a superconducting qubit coupled to the environment which is a very important approach in CQED.
@@ V.2.2 Derivation of the Bloch-Redfield master equation
$R_{abcd}=-\frac{1}{2}\sum_{\alpha\beta}\Big{\{}\delta_{bd}\sum_{n}A^{\alpha}_{an}A^{\beta}_{nc}S_{\alpha\beta}(\omega_{cn})-A^{\alpha}_{ac}A^{\beta}_{db}S_{\alpha\beta}(\omega_{ca})+\delta_{ac}\sum_{n}A^{\alpha}_{dn}A^{\beta}_{nb}S_{\alpha\beta}(\omega_{dn})-A^{\alpha}_{ac}A^{\beta}_{db}S_{\alpha\beta}(\omega_{db})\Big{\}},$ (23)
@@ V.3.2 Non-Markovian dynamics of superconducting qubits
$L_{1}\rho(t)=\gamma_{z}\big{(}(\sigma_{z}\rho(t)\sigma_{z}-\rho(t)\big{)},$ (27)
@@ V.4.2 Formalism of the Monte Carlo approach
$=\rho(t)+\delta t\Big{(}\frac{H_{\rm eff}}{i}\rho(t)-\rho(t)\frac{H^{\dagger}_{\rm eff}}{i}\Big{)}+\delta t\sum_{m}C_{m}\rho(t)C^{\dagger}_{m}+\mathcal{O}(\delta t^{2}),$ (36)
@@ V.5.1 Floquet formalism
The Floquet formalism is only appropriate to find $|\Psi(t)\rangle$ of Hamiltonian $H(t)$, if we can evaluate the Floquet states and quasienergies more easily than solving the time-dependent Schrödinger equation in Eq. (40). This formalism is a useful tool to study strongly driven periodic quantum systems. It not only considers the periodic nature of the perturbation, but it also prevents the appearance of terms that are not periodic or linear in time called secular terms Grifoni .
@@ V.5.2 Floquet master equation
In the Lindblad and Bloch-Redfield master equations, the Hamiltonian describing the system is assumed to be time-independent. Thus, these master equations should not be applied to systems with time-dependent Hamiltonians such as a driven qubit that is interacting with its environment. However, in a weak driving regime, the standard master equations still could be useful for many time-dependent problems Breuer .
@@
$S(\omega)=\int_{-\infty}^{\infty}\operatorname{d}t\exp(i\omega t){\rm Tr}_{\varepsilon}[\eta(t)\eta(0)\rho_{\varepsilon}],$ (62)
@@
$g_{k\phi}=\frac{\omega_{d}}{4\pi}\int_{0}^{2\pi/\omega_{d}}\operatorname{d}t\exp(ik\omega_{d}t){\rm Tr}_{\rm q}[\sigma c_{\phi}(t)],$ (63b)
@@ V.6 Perturbation Theory
In quantum mechanics, perturbation theory is a method to analyze the behavior of a complex quantum system by comparing it to a simpler one. The idea is to start with a simple system where we already have an exact solution of the Schrödinger equation for it, and then introduce perturbing Hamiltonian representing a weak perturbation to the system. If the perturbation is not very big, we can express the different physical values of the perturbed system such as energy levels and eigenstates as corrections to those of the simple system. These small corrections can be calculated using approximate approaches such as asymptotic series. Perturbation theory is a useful tool for explaining real quantum systems. It helps when we can’t find exact answers to the Schrödinger equation for more complicated Hamiltonians Messiah ; Sakurai .
@@ V.6.2 Calculation of decoherence rate
The Hamiltonian of the qubit $H_{\rm q}$ is that of a spin-$1/2$ particle under the influence of the static field $H_{0}=\omega_{\rm q}\hat{z}$ can be written as

### v2

@@ Abstract
Superconducting qubits are currently a leading platform for developing a scalable quantum computer. However, random and uncontrollable noises from the environment during the design and measurement of qubits lead to limitations in qubit coherence time and gate fidelity, which is a major challenge in current state-of-art for superconducting quantum computing. To advance superconducting qubits technologies, it is essential to understand and mitigate environmentally-induced errors. This requires modeling superconducting qubits as open quantum systems coupled to their surroundings. The present study aims to provide useful open quantum system approaches to analyze and quantify the interaction between superconducting qubits and their environments. We provide an accessible introduction to open quantum system formalism for newcomers to the field. For experts, we discuss recently developed methods for analyzing superconducting qubit dynamics under realistic noises. We outline how these techniques provide quantitative insights into decoherence mechanisms and how they can guide design improvements to enhance qubit coherence time. This self-contained review of open quantum system approaches can be used to model, understand, and improve superconducting qubit performance in the presence of unavoidable environmental noises.
Keywords decoherence, superconducting qubit, noise, circuit quantum electrodynamics
@@ III.1 Basic Definitions
Figure 3: A general schematic of a superconducting qubit composed of an inductor (L), a capacitor (C) and a Josephson junction with critical current (Ic). Source Kwon .
@@
in which the ground states of the qubit are eigenstates of charge. It means that they are only dependent on the number of excess Cooper pairs in a disconnected superconducting island. Single CPB and transmon are well-known examples. The transmon significantly improved upon earlier charge qubits by operating in a regime with a much larger ratio of Josephson energy to charging energy. This design choice exponentially suppresses the qubit’s sensitivity to charge noise, a major source of dephasing in earlier CPBs, leading to substantially longer coherence times, albeit at the cost of reduced anharmonicity. Moreover, it shows well-developed control and readout schemes using microwave signals.
@@
consist of a superconducting loop interrupted by one or more Josephson junctions. Their quantum states are typically superpositions of states corresponding to persistent superconducting currents flowing in opposite directions around the loop, creating a magnetic flux. Flux qubits normally exhibit large anharmonicity, comparable to or greater than early CPB charge qubits. Compared to transmon qubits, which sacrifice some anharmonicity for enhanced charge noise insensitivity, flux qubits can offer a different balance of these properties and often achieve high coherence. C-shunted flux qubits and fluxonium qubits are examples. Unlike the transmon, the fluxonium typically incorporates a junction array or a large superinductor, shunting a small Josephson junction. This design creates a different energy spectrum, characterized by a small plasma frequency and strong anharmonicity. Fluxonium qubits can exhibit excellent coherence times, particularly robust against flux noise at specific operating points, and offer a rich parameter space for engineering qubit properties. However, it requires complex fabrication (superinductor is large and made of many junctions) and more complicated control/readout than transmon due to a more complex energy spectrum.
@@
which have bigger Josephson junctions compared to charge qubits. Their states relate to two levels of current energy in a Josephson junction which two levels are defined by quantum oscillations of the phase difference between the two sides of the junction. In phase qubits, the different states are usually determined by the direction of a flowing current in the Josephson loop. Phase qubits were among the early successful demonstrations of coherent quantum phenomena in macroscopic superconducting circuits. Moreover, experiments with phase qubits in the early 2000s (e.g., by John Martinis’ group) were pivotal in demonstrating single-qubit operations (initialization, Rabi oscillations, readout) and contributed significantly to establishing superconducting circuits as a viable platform for quantum computing Devoret .
@@ III.2 Hamiltonian
In this section, we define a foundational Hamiltonian form, given in FIG. 3, that describes many common superconducting qubits based on a single Josephson junction with critical current (Ic) along with linear inductive (L) and capacitive (C) elements. Note that this Hamiltonian serves as a starting point for understanding various qubit types such as charge, flux, and transmon qubits, though more complex or specialized designs might necessitate extended or different Hamiltonians Kwon ; Vool .
@@ III.3.1 Formalism of decoherence
The formalism of decoherence in superconducting qubits is similar to other quantum systems and includes two main steps: relaxation and dephasing Dong ; Feng . These mechanisms lead to the exponential vanishing of off-diagonal elements of the density matrix $\rho$ of the qubit which shows the overlap between two states of a qubit.
@@ III.3.2 Source of decoherence
is a common noise in solid-state devices. This noise originates fluctuating charges in the structural disorder or charge traps in the dielectrics and tunnel barrier of the junction. It can be modeled as an ensemble of fluctuating two-level systems or as bulk dielectric loss Wang ; Dial . To mitigate the effects of charge noise, we can use the etching and passivation to remove amorphous layers, crystalline dielectrics and ordered Josephson junction barriers Siddiqi .
@@ IV Circuit Quantum Electrodynamics (CQED)
CQED emerged as a highly influential research field approximately two decades ago, rapidly becoming a cornerstone for superconducting quantum computing Blais2 ; Wallraff ; Girvin . CQED uses theoretical and experimental methods to manipulate quantum states for technological applications. Currently, it plays an essential role in quantum information processing based on superconducting qubits Blais . Moreover, CQED is now a leading strategy in single photon detectors Besse ; Kono ; Lescanne , development of quantum-limited amplifiers Clerk1 ; Roy , hybrid quantum systems Clerk2 and semiconducting quantum dots Burkard .
@@
$H=\hbar\omega_{\rm c}(a^{\dagger}a+\frac{1}{2})+\frac{\hbar\omega_{\rm q}}{2}\sigma_{z}+\hbar g(a^{\dagger}\sigma^{-}+\sigma^{+}a)+H_{\rm c\varepsilon}+H_{\rm q\varepsilon}.$ (6)
The first term represents cavity harmonic oscillator energy that $\omega_{\rm c}$ is the angular oscillation frequency of the relevant cavity mode. The second term shows the qubit energy in which the qubit frequency $\omega_{\rm q}$ is the energy difference between the ground state $|g\rangle$ and the excited state $|e\rangle$ of the qubit. The third term denotes the coupling between qubit and cavity with the strength of interaction $g$ in which the rotating wave approximation (RWA) is applied Knight ; Walls . This term shows the coherent exchange of excitation between the qubit and cavity and describes a dipole interaction where an atom can absorb ($\sigma^{+}a$) and emit ($a^{\dagger}\sigma^{-}$) a photon from/to the field at rate $g$. The two last terms show the coupling of cavity and qubit with the environment, respectively, that causes relaxation and dephasing of the qubit. Moreover, $a$ and $a^{\dagger}$ are annihilation and creation operators of the resonator photons, respectively and $\sigma_{z}=|e\rangle\langle e|-|g\rangle\langle g|$ is the atomic inversion operator and the operators $\sigma^{-}=|g\rangle\langle e|$ and $\sigma^{+}=|e\rangle\langle g|$ are the raising and lowering operators of the qubit. The detuning of a qubit from the cavity can be defined as $\omega_{\rm cq}=\omega_{\rm q}-\omega_{\rm c}$.
@@ V.1.1 Non-unitary evolution
To describe the state of an open quantum system as a mixed state, we use the density matrix formalism. A density matrix $\rho$ is a Hermitian operator describes a probability distribution of quantum states $\left|\psi_{n}\right>$ in a matrix representation that defined as $\rho=\sum_{n}p_{n}\left|\psi_{n}\right>\left<\psi_{n}\right|$, where $p_{n}$ is the probability that the system is in $\left|\psi_{n}\right>$. The normalization condition of the density matrix is defined as $\operatorname{Tr}\rho=1$ where $\operatorname{Tr}$ denotes Trace operation. Moreover, for a pure quantum state, we have $\operatorname{Tr}\rho^{2}=1$ and for a mixed state $\operatorname{Tr}\rho^{2}<1$. In general, the time evolution of a density matrix $\rho$ is the basis of the master equations in the theory of open quantum systems Schlosshauer1 ; Breuer .
@@ V.1.2 Derivation of the Lindblad master equation
$\frac{\operatorname{d}\rho_{\rm q}(t)}{\operatorname{d}t}=-i[H_{\rm q}(t),\rho_{\rm q}(t)]+\Gamma_{1}\Big{(}\sigma_{-}\rho_{\rm q}\sigma_{+}-\frac{1}{2}\{\sigma_{-}\sigma_{+},\rho_{\rm q}\}\Big{)}+\Gamma_{\phi}\Big{(}\sigma_{z}\rho_{\rm q}\sigma_{z}-\rho_{\rm q}\Big{)}.$ (15)
@@
Note that Eq. (15) presents a general phenomenological form of the Lindblad master equation for a single qubit experiencing energy relaxation and pure dephasing. More specific forms of the Lindblad equation are often derived for particular physical setups, like a qubit coupled to a cavity resonator which then interacts with a thermal bath. In such cases, the effective qubit relaxation rate can be expressed in terms of system parameters like the qubit-cavity coupling strength, cavity decay rate, and detuning Blais3 .
@@ V.2.1 Framework
The evolution of a system and its environment is taken from the total Hamiltonian and traces over the degrees of freedom of the environment. Although various methods have been introduced in the approach of perturbative analysis of the interaction between the system and the environment such as Bloch-Redfield Bloch ; Redfield , Nakajima-Zwanzig Nakajima ; Zwanzig , Schoeller-Schon Schoeller . Under certain approximations, such as the weak coupling limit and the Markov approximation, the Nakajima-Zwanzig formalism can be shown to lead to a master equation of the Redfield form. The relationships between these various perturbative approaches are nuanced, often depending on the specific approximations made in their derivation.
@@
The Bloch-Redfield master equation has been, and continues to be, used extensively to study the dissipative behavior of qubits interacting with their environment.Whitney . In the following, we analyze the formalism to derive the Bloch-Redfield master equation for a superconducting qubit coupled to the environment which is a very important approach in CQED.
@@ V.2.2 Derivation of the Bloch-Redfield master equation
$R_{abcd}=-\frac{1}{2}\sum_{\alpha\beta}\Big{\{}\delta_{bd}\sum_{n}A^{\alpha}_{an}A^{\beta}_{nc}S_{\alpha\beta}(\omega_{cn})-A^{\alpha}_{ac}A^{\beta}_{db}S_{\alpha\beta}(\omega_{ca})+\delta_{ac}\sum_{n}A^{\alpha}_{dn}A^{\beta}_{nb}S_{\alpha\beta}(\omega_{dn})-A^{\alpha}_{ac}A^{\beta}_{db}S_{\alpha\beta}(\omega_{db})\Big{\}}.$ (23)
For instance, when modeling a transmon qubit where 1/f flux noise is a dominant dephasing mechanism at its optimal point, one might use the Bloch-Redfield formalism with an appropriate flux noise spectral density as input to Eq. (23).
@@
While the Bloch-Redfield equation offers a more direct connection to the microscopic environment properties than the phenomenological Lindblad approach, care must be taken as it does not always guarantee positivity, unlike the Lindblad form which can be achieved via a secular approximation. Moreover, while Markovian master equations like Lindblad and Bloch-Redfield are widely applicable, situations involving strong system-environment coupling or structured environments with memory necessitate non-Markovian approaches, which we discuss next.
@@ V.3.2 Non-Markovian dynamics of superconducting qubits
$L_{1}\rho(t)=\gamma_{z}\big{(}(\sigma_{z}\rho(t)\sigma_{z}-\rho(t)\big{)}.$ (27)
@@
In the next section, we use an approach based on an effective Hamiltonian to describe the non-Hermitian dynamics of open quantum systems.
@@ V.4.2 Formalism of the Monte Carlo approach
$=\rho(t)+\delta t\Big{(}\frac{H_{\rm eff}}{i}\rho(t)-\rho(t)\frac{H^{\dagger}_{\rm eff}}{i}\Big{)}+\delta t\sum_{m}C_{m}\rho(t)C^{\dagger}_{m}+\mathcal{O}(\delta t^{2}).$ (36)
@@ V.5.1 Floquet formalism
The Floquet formalism is only appropriate to find $|\Psi(t)\rangle$ of Hamiltonian $H(t)$, if we can evaluate the Floquet states and quasienergies more easily than solving the time-dependent Schrödinger equation in Eq. (40). This formalism is a useful tool to study strongly driven periodic quantum systems. This non-perturbative formalism is a useful tool to study strongly driven periodic quantum systems. It explicitly handles the time-periodic nature of the driving Hamiltonian (which can be thought of as a time-dependent perturbation to the undriven system’s dynamics), and by transforming to the Floquet basis, it helps in avoiding secular terms that can arise in naive perturbative treatments of periodically driven systems Grifoni .
@@ V.5.2 Floquet master equation
$S(\omega)=\int_{-\infty}^{\infty}\operatorname{d}t\exp(i\omega t){\rm Tr}_{\varepsilon}[\eta(t)\eta(0)\rho_{\varepsilon}].$ (62)
@@
$g_{k\phi}=\frac{\omega_{d}}{4\pi}\int_{0}^{2\pi/\omega_{d}}\operatorname{d}t\exp(ik\omega_{d}t){\rm Tr}_{\rm q}[\sigma c_{\phi}(t)].$ (63b)
@@
As we mentioned before, the Floquet theory is a non-perturbative approach to describe the effect of an oscillating field on the superconducting qubits. In the next section, we use a perturbative method to analyze the dynamics of the superconducting qubits as open quantum systems.
@@ V.6 Perturbation Theory
In quantum mechanics, perturbation theory is a method to analyze the behavior of a complex quantum system by comparing it to a simpler one. The idea is to start with a simple system where we already have an exact solution of the Schrödinger equation for it, and then introduce a perturbing Hamiltonian representing a weak perturbation to the system. If the perturbation is not very big, we can express the different physical values of the perturbed system such as energy levels and eigenstates as corrections to those of the simple system. These small corrections can be calculated using approximate approaches such as asymptotic series. Perturbation theory is a useful tool for explaining real quantum systems. It helps when we can’t find exact answers to the Schrödinger equation for more complicated Hamiltonians Messiah ; Sakurai .
@@ V.6.2 Calculation of decoherence rate
The Hamiltonian of the qubit such as quantronium $H_{\rm q}$ is that of a spin-$1/2$ particle under the influence of the static field $H_{0}=\omega_{\rm q}\hat{z}$ can be written as
@@
So far, we haven’t investigated the interaction between a superconducting qubit and an electromagnetic field through the transmission line. In the next section, we use input-output theory to obtain information about the fields radiated by the superconducting qubits that is of practical importance in qubit readout.